One day a store sold a total of 178 pencils and notebooks. The cost of a pencil is $0.25 and the cost of a notebook is $0.50. The store collected $58.75 from the sale of these items. The store lost the receipt for the day. How would you help the store determine how many of each item the store sold that day?
Please help me solve this and please tell me the "method" you would use to solve this. Along with that, please explain why you would choose that method.
Thank you, and please respond with haste.

Answer by fcabanski(1391)

Call pencils P and notebooks N.


P + N = 178, which is the total number of items sold.


.25P + .5N = 58.75 - the price per pencil times the number of pencils sold, plus the price per notebook times the number of notebooks sold, equals the total sales amount.


P + N = 178 so P = 178-N. Substitute that into the second equation.


.25(178-N) +.5N = 58.75


44.50 - .25N + .5N = 58.75


.25N = 14.25


N = 14.25/.25 = 57


P = 178 - 57 = 121


Check: .25*121 + .50*57 = 30.25+28.50 = 58.75